OpenStudy (anonymous):

x + y = 90 degrees; tan(x) = 14.504 What is tan(y)?

OpenStudy (anonymous):

Get The value of x from Tan(x)^-1=14.504 then substitute with that value in the equation x+y=90 ,hence you get the value of y subsitiute in tan(y)

OpenStudy (anonymous):

Arctan of 14.504 = x

OpenStudy (anonymous):

I get that tan(y) = 3.865599773 rounding to 3.866 when I enter the answer, it is wrong.

OpenStudy (anonymous):

How did you get tan(y) = 3.866?

OpenStudy (anonymous):

well. I did 90 - 14.504 = 75.496. Next I did tan(75.496) = 3.866

OpenStudy (anonymous):

But x doesn't equal 14.504, tan(x) does. Using the acrtangent function tan^-1 (14.504) = x

OpenStudy (anonymous):

okay. I get 86.05589836

OpenStudy (anonymous):

so, does y= 90-86.05589836?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so y = 3.94410164 then do I take the tan of that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you very much. I appreciate your help.

OpenStudy (anonymous):

Your welcome.